# A Novel Lexicon for the Moral Foundation of Liberty

### v1

@@ 4.3 Evaluation
. We analyze the in-domain performance of our lexicons by testing them on a left-out set of the datasets they are generated from (consisting of 20% of the original data). Notice how, depending on the dataset, the task is slightly different due to the different type of annotations: (i) for the BLM and Elect. datasets, the classifier should predict whether the document expresses notions of liberty/oppression or is neutral; (ii) for the Reddit and WikiCon datasets, it should predict whether a document expresses the libertarian or the conservative point of view..
To avoid overfitting, each model is training on the training set of the respective balanced dataset (see Table 1), leaving the test set for evaluation.
Out-of-domain evaluation
@@
. We perform a series of out-of-domain experiments, testing how well the lexicons can generalize to different domains. In particular, we measure the performance of: (i) the lexicons generated from Reddit and WikiCon used on the BLM and Election datasets; (ii) the lexicons generated from BLM and Elect. used on the Reddit and WikiCon datasets; (iii) the lexicons generated from BLM and Elect., trained on BLM/Elect. and tested on the Vaccine dataset.
@@
Here the different annotation schemes and domains could potentially pose a bigger challenge for the classifier; however, while in the first two cases we can expect to see the impact of the different vocabulary, the train/test split is still coming from the same dataset (in other words, the model has to learn a task using sub-optimal features, but having “coherent” data for training and testing). The Vaccine dataset is instead used to test the out-of-domain performance of the lexicons when the annotation scheme of the evaluation dataset (i.e., the presence or absence of liberty foundation) is coherent with the annotation scheme of the dataset from which the lexicons are generated (BLM and Elect.), but no ideal training data is available (as we are training on an annotated dataset -BLM or Elect- that is different from the Vaccine test dataset).
For all experiments we utilized logistic regression (Alpaydin, 2020) and represented each document using a vector of the same length as each lexicon vocabulary. The feature vectors are constructed as follows; each document is represented by a vector of equal size to the lexicon. For those tokens in the document present in the lexicon, the vector contains the respective polarity score, otherwise zero. Since this type of representation dramatically simplifies the linguistic information present in the document, we enhance the classification design with two more experiments. We extend each vector representation with the “statistical summary” functions, namely the average, maximum, median, variance, and a peak-to-peak score of the lexicon values of that document. This offers the learning models a more complete view of the text.
Figure 1: Diagram of aggregation through combined representation for an example case with 4 individual lexicons.
Combined lexicons
@@
. To test whether it is possible to obtain a more “general purpose model”, we evaluate two ways of combining the information coming from the different lexicons: (i) the lexicon overlap described in Section 4.2, which averages the values for words that appear in multiple lexicons; (ii) the combined representation, which is not a lexicon, but a method of learning a unified representation by taking into account all available lexicons.
The advantage of the first method is its simplicity, and that it results in an interpretable lexicon. On the other hand, the combined representation allows a learning model to observe simultaneously all information contained in the individual lexicons (including words not shared among them); the model might then be able to exploit existing interactions among them.
@@
While the strength of this second approach is that it provides a comprehensive representation obtained through all individual lexicons, overfitting may occur given the large dimensionality of representation. To avoid such issue, we include a feature selection mechanism in the learning model so that the dimension of the feature vector can be reduced. Our approach is based on the Singular Value Decomposition technique (SVD) (Halko et al., 2011) for transforming the representation of a single lexicon into a continuous vector, which is then input to a machine learning algorithm, in our case a logistic regressor. Figure 1 illustrates the hierarchical structure of the proposed model.
Since these two approaches take into account all generated lexicons, we can consider the results of the lexicon overlap and combined representation methods: (i) an in-domain evaluation when applied to the BLM, Elect., Reddit or WikiCon datasets (since they are used to generate the individual lexicons), (ii) an out-of-domain evaluation, when applied to the Vaccine dataset (since this is only used as a test set).
@@
As baseline models, we train two classifiers using a unigram representation that includes a frequency-selected vocabulary of sizes 1,000 and 10,000 tokens respectively. The two sizes are comparable to the size of the obtained lexicons (see Table 2).
@@
To assess the general quality of the lexicons and obtain an overall ranking of their performance, we have performed the Friedman statistical test over all the evaluation results (Araque et al., 2017; Demšar, 2006). In the Friedman test a lower ranking implies a better result for a certain method in comparison to the rest. In case of ties, these are resolved by averaging the obtained ranks. The Friedman test has been performed with $\alpha=0.05$, rejecting the null hypothesis. We report the macro-averaged F-score as well as the Friedman rank for the overall evaluation of each resource.
@@ 5.1 Lexicon Generation
Table 2 shows the vocabulary size for all the generated lexicons. As previously mentioned, we generated one lexicon from each of the datasets (see Table 1) using the two proposed methods WE and CS, except for the Vaccine dataset, which is used only for the out of domain evaluation. Finally, using the overlapping approach, we obtained a representation that combines the shared tokens from the individual lexicons, using their average scores related to liberty. For the lexicons generated with the Compositional Semantics method, we applied a 10 frequency cut-off for Reddit and WikiCon. Due to the limited number of annotated instances in the BLM and Election datasets, we have set a 6 and 3 frequency cut-off, respectively. These cut-off variations have been experimentally validated on the training data, and are in line with the literature Araque et al. (2019). For the lexicons generated with the WE method, the same frequency cut-off has been applied to the Reddit and WikiCon lexicons, while we did not apply any cut-off for the BLM and Election lexicons generated this way, to increase their vocabulary size. Generally, we have observed that the two methods show a dependency between the number of annotated instances in the training data and the resulting vocabulary size.
For the overlapping lexicons and due to space limitations, we report the data for the lexicon generated using as selection (cut-off) parameter 40%, which resulted the best combination in the supervised evaluation. To do this, we evaluated the selection parameter in the range [10%, 20%, …, 100%] on the train sets of the considered datasets using 10-fold cross-validation. This selection justifies the more limited number of tokens with respect to the aggregation of all the lexicons’ tokens.
@@ 5.2 Lexicon Evaluation
Unigram (1000) 50.90 50.12 66.80 83.84 52.11 43.66 9.1
Unigram (10,000) 51.84 51.81 68.39 88.10 51.17 49.70 8.8
BLM (CS) 51.53 48.38 61.99 81.43 47.12 33.33 8.6
Elect. (CS) 51.21 50.52 63.59 84.89 41.44 35.67 10.8
@@
WikiCon (CS) 50.24 49.65 68.12 90.34 52.63 42.10 5.8
BLM (WE) 46.20 53.23 65.37 85.56 52.06 57.78 7.8
Elect. (WE) 49.76 47.50 66.84 88.68 49.15 58.15 5.6
Reddit (WE) 51.19 55.32 64.32 88.50 53.72 54.85 8.0
WikiCon (WE) 50.67 53.58 65.84 88.64 51.86 54.09 6.8
Lexicon Overlap 52.40 50.63 67.54 89.29 52.44 42.91 3.5
@@
We obtain further confirmation for this hypothesis by looking at the Friedman test: when considering all evaluation combinations, the Combined Representations ranks as the best approach, followed by the lexicon overlap. This is to be expected, as this method makes use of all lexicons simultaneously, and learns internal representations that can be exploited by a machine learning model.
@@
Domain Specific Insights
A recurrent pattern is that the overlapping lexicon outperforms the other lexicons in both in-domain and out-of-domain experiments. To gain more insights on the effect of the social context on the moral nuances a specific lemma may have, we employed the TOMEA approach proposed by Liscio et al. (2023). According to TOMEA, the overlapping lexicon differs the most with respect to the BLM lexicons generated by both the CS and the WE methods with scores .17 and .13, respectively. Glancing into the most distant words, we have “fake”, “lawmaker”, “elected”, “supporters”, “antifa”, “openly”, “tweets”, “sympathizer”, “tyranny”, “globalist”, “dead”, to be considered more oppressive than the average lexicons in the individual BLM lexicons than in the others. Such domain specific nuances may be important when analysing a specific argument but can introduce biased in the models when analysing broader subjects. TOMEA to this respect is a valuable method to gain insights and foster the transparency and accountability of the findings.
@@ 6 Conclusion
The aim of this study was to provide a lexical resource for the moral foundation of liberty able to generalise across various domains. The “Liberty/Oppression” foundation expresses people’s inclinations towards autonomy and their resistance to dominion.
Our contributions to the current state of the art are manifold. Firstly, we provide an enriched version of the MFTC annotations for the BLM and ELECT datasets as per the liberty foundation. Further, we generated a series of lexicons with two complementary approaches and thoroughly evaluated them via a series of both in- and out-of domain experimental scenarios. Aside the individual lexicons, we also proposed a final version combining the information from both approaches which is the resource we propose as the final Liberty Moral Lexicon. This resource showed solid generalisability potentials. Moreover, we design a combined representation that exploits information in all generated lexicons, thus offering a classifier a more comprehensive representation. As seen in the experimental evaluation, the generated resources capture relevant knowledge that can be leveraged to assessing liberty in texts. These insights are supported by the Friedman test, that offers us a ranking of methods. Finally, by employing the TOMEA method, we provide insights into the dynamics of linguistic variability according to the context in which the word is used.

### v2

@@ 4.3 Evaluation
We analyze the in-domain performance of our lexicons by testing them on a left-out set of the datasets they are generated from (consisting of 20% of the original data). Notice how, depending on the dataset, the task is slightly different due to the different type of annotations: (i) for the BLM and Elect. datasets, the classifier should predict whether the document expresses notions of liberty/oppression or is neutral; (ii) for the Reddit and WikiCon datasets, it should predict whether a document expresses the libertarian or the conservative point of view.. To avoid overfitting, each model is training on the training set of the respective balanced dataset (see Table 1), leaving the test set for evaluation.
Out-of-domain evaluation.
@@
We perform a series of out-of-domain experiments, testing how well the lexicons can generalize to different domains. In particular, we measure the performance of: (i) the lexicons generated from Reddit and WikiCon used on the BLM and Election datasets; (ii) the lexicons generated from BLM and Elect. used on the Reddit and WikiCon datasets; (iii) the lexicons generated from BLM and Elect., trained on BLM/Elect. and tested on the Vaccine dataset.
@@
Here the different annotation schemes and domains could potentially pose a bigger challenge for the classifier; however, while in the first two cases we can expect to see the impact of the different vocabulary, the train/test split is still coming from the same dataset (in other words, the model has to learn a task using sub-optimal features, but having “coherent” data for training and testing). The Vaccine dataset is instead used to test the out-of-domain performance of the lexicons when the annotation scheme of the evaluation dataset (i.e., the presence or absence of liberty foundation) is coherent with the annotation scheme of the dataset from which the lexicons are generated (BLM and Elect.), but no ideal training data is available (as we are training on an annotated dataset -BLM or Elect- that is different from the Vaccine test dataset). For all experiments, we used logistic regression [1] and represented each document as a vector of the same length as each lexicon vocabulary, as described in the following. Each document is represented by a vector of equal size to the lexicon. For those tokens in the document present in the lexicon, the vector contains the respective polarity score, otherwise zero. Since this type of representation dramatically simplifies the linguistic information present in the document, we enhance the classification design with two more experiments. We extend each vector representation with the “statistical summary” functions, namely the average, maximum, median, variance, and a peak-to-peak score of the lexicon values of that document. This offers the learning models a more complete view of the text.
Combined lexicons.
@@
To test whether it is possible to obtain a more “general purpose model”, we evaluate two ways of combining the information coming from the different lexicons: (i) the LML described in Section 4.2, which averages the values for words that appear in multiple lexicons; (ii) the combined representation, which is not a lexicon, but a method of learning a unified representation by taking into account all available lexicons. The advantage of the first method is its simplicity, and that it results in an interpretable lexicon. On the other hand, the combined representation allows a learning model to observe simultaneously all information contained in the individual lexicons (including words not shared among them); the model might then be able to exploit existing interactions among them.
@@
While the strength of this second approach is that it provides a comprehensive representation obtained through all individual lexicons, overfitting may occur given the large dimensionality of representation. To avoid such issue, we include a feature selection mechanism in the learning model so that the dimension of the feature vector can be reduced. Our approach is based on the Singular Value Decomposition technique (SVD) [17] for transforming the representation of a single lexicon into a continuous vector, which is then input to a machine learning algorithm, in our case a logistic regressor. Since these two approaches take into account all generated lexicons, we can consider the results of the LML and combined representation methods: (i) an in-domain evaluation when applied to the BLM, Elect., Reddit or WikiCon datasets (since they are used to generate the individual lexicons), (ii) an out-of-domain evaluation, when applied to the Vaccine dataset (since this is only used as a test set).
@@
As baseline models, we train two classifiers using a unigram representation that includes a frequency-selected vocabulary of sizes 1,000 and 10,000 tokens respectively, which are comparable to the size of the lexicons (Table 2). We benchmark LML against the only publicly-available liberty lexicon, namely the LibertyMFD [7] as an external baseline.
@@
To assess the general quality of the lexicons and obtain a ranking of their performance, we performed the Friedman statistical test over all the evaluation results [10]. In the Friedman test a lower ranking implies a better result for a certain method in comparison to the rest. In case of ties, these are resolved by averaging the obtained ranks. The Friedman test has been performed with $\alpha=0.05$, rejecting the null hypothesis. We report the macro-averaged F-score as well as the Friedman rank for the evaluation of each resource.
@@ 5.1 Lexicon Generation
Table 2 shows the vocabulary size for all the generated lexicons. As previously mentioned, we generated one lexicon from each of the datasets (see Table 1) using the two proposed methods WE and CS, except for the Vaccine dataset, which is used only for the out of domain evaluation. Finally, using the overlapping approach, we obtained a representation that combines the shared tokens from the individual lexicons, using their average scores related to liberty. For the lexicons generated with the Compositional Semantics method, we applied a 10 frequency cut-off for Reddit and WikiCon. Due to the limited number of annotated instances in the BLM and Election datasets, we have set a 6 and 3 frequency cut-off, respectively. These cut-off variations have been experimentally validated on the training data, and are in line with the literature [3]. For the lexicons generated with the WE method, the same frequency cut-off has been applied to the Reddit and WikiCon lexicons, while we did not apply any cut-off for the BLM and Election lexicons generated this way, to increase their vocabulary size. Generally, we have observed that the two methods show a dependency between the number of annotated instances in the training data and the resulting vocabulary size.
For the overlapping lexicons and due to space limitations, we report the data for the lexicon generated using as selection (cut-off) parameter 40%, which resulted in the best combination in the supervised evaluation. To do this, we evaluated the selection parameter in the range [10%, 20%, …, 100%] on the train sets of the considered datasets using 10-fold cross-validation. This selection justifies the more limited number of tokens with respect to the aggregation of all the lexicons’ tokens.
@@ 5.2 Lexicon Evaluation
Unigram (1000) 50.90 50.12 66.80 83.84 52.11 43.66 8.1
Unigram (10,000) 51.84 51.81 68.39 88.10 51.17 49.70 7.6
LibertyMFD (CS) 50.01 50.74 64.90 88.39 42.32 47.46 9.6
LibertyMFD (WE) 50.39 51.62 65.96 88.47 39.66 48.63 8.0
BLM (CS) 51.53 48.38 61.99 81.43 47.12 33.33 10.0
Elect. (CS) 51.21 50.52 63.59 84.89 41.44 35.67 12.6
@@
WikiCon (CS) 50.24 49.65 68.12 90.34 52.63 42.10 6.6
BLM (WE) 46.20 53.23 65.37 85.56 52.06 57.78 9.0
Elect. (WE) 49.76 47.50 66.84 88.68 49.15 58.15 6.0
Reddit (WE) 51.19 55.32 64.32 88.50 53.72 54.85 8.8
WikiCon (WE) 50.67 53.58 65.84 88.64 51.86 54.09 7.4
LML 52.40 50.63 67.54 89.29 52.44 42.91 4.2
@@
We obtain further confirmation for this hypothesis by looking at the Friedman test: when considering all evaluation combinations, the Combined Representations ranks as the best approach, followed by the LML. This is to be expected, as this method makes use of all lexicons simultaneously, and learns internal representations that can be exploited by a machine learning model.
@@
Domain Specific Insights.
A recurrent pattern is that the overlapping lexicon outperforms the other lexicons in both in-domain and out-of-domain experiments. To gain more insights on the effect of the social context on the moral nuances a specific lemma may have, we employed the TOMEA approach proposed in [25]. According to TOMEA, the overlapping lexicon differs the most with respect to the BLM lexicons generated by both the CS and the WE methods with scores .17 and .13, respectively. Glancing into the most distant words, we have “fake”, “lawmaker”, “elected”, “supporters”, “antifa”, “openly”, “tweets”, “sympathizer”, “tyranny”, “globalist”, “dead”, to be considered more oppressive than the average lexicons in the individual BLM lexicons than in the others. Such domain specific nuances may be important when analysing a specific argument but can introduce biased in the models when analysing broader subjects. TOMEA to this respect is a valuable method to gain insights and foster the transparency and accountability of the findings.
@@ 6 Conclusion
The aim of this study was to provide a lexical resource for the moral foundation of liberty able to generalise across various domains. The “Liberty/Oppression” foundation expresses people’s inclinations towards autonomy and their resistance to dominion. Our contributions to the current state of the art are manifold. Firstly, we provide an enriched version of the MFTC annotations for the BLM and ELECT datasets as per the liberty foundation. Further, we generated a series of lexicons with two complementary approaches and thoroughly evaluated them via a series of both in- and out-of domain experimental scenarios. Aside the individual lexicons, we also proposed a final version combining the information from both approaches which is the resource we propose as the final Liberty Moral Lexicon (LML). This combined representation exploits information in all generated lexicons, thus offering a classifier a more comprehensive representation outperforming individual expert lexicons [18]. This resource showed solid generalisability potentials. As seen in our experimental evaluation, the generated resources capture relevant knowledge that can be leveraged to assessing liberty in texts with statistical significance. Finally, by employing the TOMEA method, we provide insights into the dynamics of linguistic variability according to the context in which the word is used. This study reaffirms the critical role of lexicons in advancing moral foundation analysis by enabling context-aware evaluations of concepts like liberty. Tools like TOMEA rely on these structured resources to decode the domain-specific expressions of moral values effectively and are fundamental in fields related to communication or policymaking when accountability and transparency are a strong requirement.
Acknowledgments
We would like to thank the colleagues of the Monitoring, Indicators & Impact Evaluation Unit of the European Commission Joint Research Centre for their support. The views expressed are purely those of the authors and may not in any circumstance be regarded as stating an official position of the European Commission. OA would like to acknowledge the funding of funded by the Spanish Ministry for Economic Affairs and Digital Transformation and by the European Union - NextGenerationEU within the Programme UNICO I+D Cloud (TSI-063100-2022-0002); as well as the funding of the project CPP2023-010437 financed by the MCIN / AEI / 10.13039/501100011033 / FEDER, UE. This work has been supported by the Madrid Government (Comunidad de Madrid-Spain) under the Multiannual Agreement 2023-2026 with Universidad Politécnica de Madrid in the Line A, Emerging PhD researchers (project MORA, DOCTORES-EMERGENTES-24-9UMLXZ-37-IIGW). KK acknowledges support from the Lagrange Project of ISI Foundation funded by CRT Foundation.